# GIST: Greedy Independent Set Thresholding for Diverse Data Summarization

### v1

@@ Abstract
We propose a novel subset selection task called min-distance diverse data summarization (MDDS), which has a wide variety of applications in machine learning, e.g., data sampling and feature selection. Given a set of points in a metric space, the goal is to maximize an objective that combines the total utility of the points and a diversity term that captures the minimum distance between any pair of selected points, subject to the constraint $|S|\leq k$. For example, the points may correspond to training examples in a data sampling problem, e.g., learned embeddings of images extracted from a deep neural network. This work presents the GIST algorithm, which achieves a $\nicefrac{{2}}{{3}}$-approximation guarantee for MDDS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm. We also prove a complementary $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$. Finally, we provide an empirical study that demonstrates GIST outperforms existing methods for MDDS on synthetic data, and also for a real-world image classification experiment the studies single-shot subset selection for ImageNet.
@@ 1 Introduction
Subset selection is a challenging optimization problem with a wide variety of applications in machine learning, including feature selection, recommender systems, news aggregation, drug discovery, data summarization, and designing pretraining sets for large language models (Anil et al., 2023). Data sampling in particular is a salient problem due to unprecedented and continuous data collection. For example, LiDAR and imaging devices in one self-driving vehicle can easily capture ~80 terabytes of data per day (Kazhamiaka et al., 2021).
@@
In most subset selection tasks, we rely on the weight (or utility) of the objects to rank one over the other, and also to avoid selecting duplicate or near-duplicate objects. If we select a small subset, then we also want to ensure that the selected subset is a good representation of the original set. These utility, diversity, and coverage criteria can be expressed through objective functions, and the interesting research lies in developing efficient algorithms with strong approximation guarantees. The underlying machinery used in constrained subset selection algorithms shares many similarities with techniques from other areas of combinatorial optimization such as submodular maximization, $k$-center clustering, and convex hull approximations.
In this work, we study the problem of selecting a set of points in a metric space that maximizes an objective that combines their utility and a minimum pairwise-distance diversity measure. Concretely, given a set of points $V$, weight function $w:V\rightarrow\mathbb{R}_{\geq 0}$, and distance function $\textnormal{dist}:V\times V\rightarrow\mathbb{R}_{\geq 0}$, our goal is to solve the cardinality-constrained optimization problem:
$S^{*}$ $=\operatorname*{arg\,max}_{S\subseteq V}~{}\frac{1}{k}\sum_{v\in S}w(v)+\texttt{div}(S)$
subject $\hskip 13.6572pt\text{subject to}~{}~{}|S|\leq k$
@@
where the diversity term $\texttt{div}(S)\coloneqq\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ encourages maximizing the minimum distance between all pairs of distinct points in the selected subset. Bhaskara et al. (2016) call $\texttt{div}(S)$ the min-min diversity objective and Meinl et al. (2011) call it $p$-dispersion maximization, where it is known to admit a $\nicefrac{{1}}{{2}}$-approximation ratio subject to the strict cardinality constraint $|S|=k$.
@@
Other diversity objectives such as the sum-of-min-distances $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ and sum-of-all-distances $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ have also been studied (Meinl et al., 2011; Bhaskara et al., 2016). While previous works have looked at maximizing the diversity objective $\texttt{div}(S)$, our formulation that introduces the utility term has not been studied before. In particular, Bhaskara et al. (2016) observe that min-min is a powerful diversity objective, but point out a major challenge: it is highly non-monotonic, i.e., it can exhibit a stark decrease in the objective value by adding a single new point. Our work shows that the use of a utility term along with the min-min can alleviate some of these pitfalls, and allows us to inherit the strong benefits of this strict diversity objective.
@@ 1.1 Our contributions
  • We propose a novel subset selection problem called min-distance diverse data summarization (MDDS), where the goal is to maximize an objective function that combines a linear weight term with the min-min distance diversity term $\texttt{div}(S)$ subject to the cardinality constraint $|S|\leq k$. This problem has a wide variety of applications in machine learning, e.g., data denoising and active learning.
  • We give a $\nicefrac{{2}}{{3}}$-approximation algorithm for MDDS called GIST that approximates a series of maximum independent set problems on geometric intersection graphs with a bicriteria greedy algorithm and returns the best solution.
  • We prove a complementary $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$, for general metric spaces by reducing from approximate maximum clique. We also prove that MDDS is APX-complete for Euclidean metrics.
  • Our experiments show that GIST overcomes the shortcomings of existing methods for MDDS on synthetic data. We also demonstrate how GIST can be used to create better single-shot subsets of training data for an image classification benchmark on ImageNet (Russakovsky et al., 2015) compared to margin sampling and $k$-center algorithms.
1.2 Background and related work
@@
A closely related problem is $k$-center, which aims to find a subset of $k$ centers that minimize the largest distance of any point to its closest center in $S$. Concretely, the objective is $\max_{v\in V}\min_{c\in S}\textnormal{dist}(v,c)$, and a greedy $2$-approximation algorithm is widely used in active learning applications (Sener and Savarese, 2018). The main difference between our objective function and $k$-center is the ability to benefit from the utility term, e.g., uncertainty values in an active learning setting. The second difference is that we relax the need to cover every point in the original dataset, which is particularly beneficial for large datasets, where we prefer to avoid instability due to certain corner or outlier points.
Bhaskara et al. (2016) study various distance-based diversity metrics subject to a cardinality constraint. In particular, they give a $\nicefrac{{1}}{{8}}$-approximation algorithm for the sum-min diversity objective defined as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ by rounding a linear programming relaxation, and they complement this result by proving a conditional $\nicefrac{{1}}{{2}}$-hardness of approximation under the planted clique assumption. They also study the min-min objective $\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$, which is the same as $\texttt{div}(S)$. They do not have a linear weight function in their objective, so they enforce a strict cardinality constraint $|S|=k$; otherwise, it is always better to output at most one or two points. Meinl et al. (2011) study the same objective subject to $|S|=k$ under the name $p$-dispersion (also called max-min or remote-edge) diversity maximization. This cardinality issue does not arise in our formulation since we need to optimize the trade-off between competing linear weight (utility) and diversity terms.
Moumoulidou et al. (2020), Addanki et al. (2022), and Mahabadi and Trajanovski (2023) initiated the study of diversity maximization under fairness constraints. Given a partitioning of the dataset, they ensure that a prespecified number of points $k_{i}$ are selected from each group $i\in[m]$. This fairness constraint considerably differentiates their problem setting from ours (i.e., $|S|=k=\sum_{i=1}^{m}k_{i}$). Moumoulidou et al. (2020); Addanki et al. (2022) maximize the $\texttt{div}(S)$ objective, and Mahabadi and Trajanovski (2023) consider the sum-of-pairwise distances $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ and sum-min objectives. None of these works, however, add a linear weight function to the objective.
@@
In the realm of dataset curation and active learning, there have been several lines of work that study the combination of utility and diversity terms. Ash et al. (2020) invoke $k$-means++ seeding over the gradient space of a model in order to balance uncertainty and diversity. Wei et al. (2015) introduce several submodular objectives, e.g., facility location, and use them to diversify a set of uncertain examples in each active learning iteration. Citovsky et al. (2021) cluster examples which are represented by embeddings extracted from the penultimate layer of a partially trained DNN, and then use these clusters to diversify uncertain examples in each iteration. Our work differs from these and several others (e.g., Kirsch et al. (2019); Zhdanov (2019)) in that we directly incorporate both the utility and diversity terms in our objective function.
@@ 2 Preliminaries
We are given a set of $n$ points $V\subseteq M$ in the metric space $(M,\textnormal{dist})$, where $\textnormal{dist}:M\times M\rightarrow\mathbb{R}_{\geq 0}$ is the distance function. We are also given weights $w:V\rightarrow\mathbb{R}_{\geq 0}$ representing the utility of each point. We extend these functions to take subsets $S\subseteq V$ as input in the standard way: $w(S)=\sum_{v\in S}w(v)$ and $\textnormal{dist}(u,S)=\min_{v\in S}\textnormal{dist}(u,v)$. For the sake of completeness, we define $\textnormal{dist}(u,\varnothing)=\infty$.
Intersection graph.
Let $G_{d}(V)$ denote the intersection graph of $V$ for distance threshold $d$, i.e., the graph with node set $V$ and edge set
$E=\{(u,v)\in V^{2}:\textnormal{dist}(u,v)<d\}.$
This graph is well-defined for any metric space and distance $d\in\mathbb{R}$. Note that for any independent set $S$ of $G_{d}(V)$ and pair of distinct nodes $u,v\in S$, we have $\textnormal{dist}(u,v)\geq d$.
Definition 2.1.
This work focuses on the min-distance diversity reward function
@@
This reward function encourages selecting a well-separated set of points (generalizing the notion of data deduplication). To make $\texttt{div}(S)$ well-defined if $S$ is empty or consists of a singleton point, we set the diversity to be the diameter of the whole ground set $V$, implying that it is always beneficial to select at least two points to maximize the value of the selected set.
For any $\alpha\in[0,1]$, we define the objective function
$f(S)=\alpha\frac{1}{k}\sum_{v\in S}w(v)+(1-\alpha)\texttt{div}(S),$ (1)
which allows us to control the trade-off between the average utility and the diversity of set $S$. The goal of the min-distance diverse data summarization problem (MDDS) is to maximize $f(S)$ subject to a cardinality constraint $k$:
@@
Simple $\nicefrac{{1}}{{2}}$-approximation algorithm.
Equation 1 has two parts, each of which is easy to optimize in isolation. To maximize the $w(S)$ term, we simply select the $k$ heaviest points. In contrast, to maximize the $\texttt{div}(S)$ term, we select two points whose distance is the diameter of $V$. Outputting a subset that corresponds to the maximum of these two terms gives a $\nicefrac{{1}}{{2}}$-approximation guarantee. To see this, observe that
$\textnormal{ALG}_{\textnormal{simple}}$ $=\max\left\{\alpha\frac{1}{k}\max_{|S|\leq k}w(S),(1-\alpha)\max_{u,v\in V}\textnormal{dist}(u,v)\right\}$
$\geq\frac{1}{2}\cdot\alpha\frac{1}{k}\max_{|S|\leq k}w(S)+\frac{1}{2}\cdot(1-\alpha)\max_{u,v\in V}\textnormal{dist}(u,v)$
$\geq\frac{1}{2}\cdot\left(\alpha\frac{1}{k}w(S^{*})+(1-\alpha)\texttt{div}(S^{*})\right)=\frac{1}{2}\cdot\textnormal{OPT},$
where $\texttt{div}(S)$ is defined in Definition 2.1. Our main goals are to improve over this simple baseline and prove complementary hardness results.
@@ 3 Algorithm
We now present the GIST algorithm, which achieves a $(\nicefrac{{2}}{{3}}-\varepsilon)$-approximation for the MDDS problem. It works by repeatedly calling the GreedyIndependentSet heuristic to find an independent set of the intersection graph $G_{d}(V)$ for various distance thresholds $d$. The subroutine GreedyIndependentSet sorts the points by their weight in a non-increasing order breaking ties arbitrarily. It starts with an empty set of selected points, and then it scans the points according to the sorted order. For each point, GreedyIndependentSet selects the point as long as its distance to the set of selected points so far is at least $d$ and the number of selected points does not exceed $k$.
@@
Algorithm 1 Diversified data summarization using greedy-weighted maximum independent sets.
1:function GIST(points $V$, weights $w:V\rightarrow\mathbb{R}_{\geq 0}$, budget $k$, error $\varepsilon$)
2:     Initialize $S\leftarrow\textnormal{{GreedyIndependentSet}}(V,w,0,k)$triangleright $k$ heaviest points
@@
9:         Set $T\leftarrow\textnormal{{GreedyIndependentSet}}(V,w,d,k)$
@@
1:function GreedyIndependentSet(points $V$, weights $w:V\rightarrow\mathbb{R}_{\geq 0}$, distance $d$, budget $k$)
@@
3:     for $v\in V$ sorted by non-increasing weight do
4:         if $\textnormal{dist}(v,S)\geq d$ and $|S|<k$ then
5:              Update $S\leftarrow S\cup\{v\}$               return $S$
@@
For any $\varepsilon>0$, GIST outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{2}}{{3}}-\varepsilon)\cdot\textnormal{OPT}$.
The main building block in our design and analysis of GIST is the following lemma, which is inspired by the greedy $2$-approximation algorithm for metric $k$-centers.

### v2

@@ Abstract
We introduce a novel subset selection problem called min-distance diversification with monotone submodular utility (MDMS), which has a wide variety of applications in machine learning, e.g., data sampling and feature selection. Given a set of points in a metric space, the goal of MDMS is to maximize an objective function combining a monotone submodular utility term and a min-distance diversity term between any pair of selected points, subject to a cardinality constraint. We propose the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm. We also prove that it is NP-hard to approximate to within a factor of $0.5584$. Finally, we demonstrate that GIST outperforms existing benchmarks for on a real-world image classification task that studies single-shot subset selection for ImageNet.
@@ 1 Introduction
Subset selection is a challenging combinatorial optimization problem with a wide variety of applications in machine learning, including feature selection, recommender systems, drug discovery, data summarization, and designing pretraining sets for large language models (Team et al., 2024). Data sampling in particular is a salient problem due to unprecedented and continuous streams of data collection. For example, LiDAR and imaging devices in one self-driving vehicle can easily capture ~80 terabytes of data per day (Kazhamiaka et al., 2021).
@@
In many subset selection tasks, we rely on the utility (or weight) of an object to rank one over the other, and also to avoid selecting duplicate or near-duplicate objects. If we select a small subset, then we must ensure that the selected set is a good representation of the original set. These utility, diversity, and coverage criteria are expressed through objective functions, and the interesting research lies in developing efficient algorithms with strong approximation guarantees. The algorithmic tools used in constrained subset selection shares many similarities with techniques from other areas of combinatorial optimization such as submodular maximization, $k$-center clustering, and convex hull approximations.
This work introduces a novel subset selection problem called min-distance diversification with monotone submodular utility (MDMS). Given $n$ points $V$ in a metric space, a monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$, and a cardinality constraint $k$, the goal is to solve
$S^{*}$ $=\operatorname*{arg\,max}_{S\subseteq V}~{}g(S)+\texttt{div}(S)$ (1)
subject $\hskip 13.6572pt\text{subject to}~{}~{}|S|\leq k,$
@@
where the diversity term $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ encourages maximizing the min distance between any two distinct points in the selected set. Meinl et al. (2011) call $\texttt{div}(S)$ the $p$-dispersion measure, and Bhaskara et al. (2016) call it the min-min diversity objective, where is known to admit a $\nicefrac{{1}}{{2}}$-approximation ratio subject to the strict cardinality constraint $|S|=k$.
@@
Other diversity terms such as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ and $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ have also been studied (Meinl et al., 2011; Bhaskara et al., 2016; Borodin et al., 2017). While previous works have explored maximizing $\texttt{div}(S)$ in isolation, MDMS introduces a utility term and has not been studied before. Bhaskara et al. (2016) remark that $\texttt{div}(S)$ is a powerful diversity objective, but point out a major challenge: it is highly non-monotonic, i.e., it can exhibit a stark decrease in the objective value by adding a single new point. Our work shows combining a monotone submodular utility term with $\texttt{div}(S)$ mitigates some of the challenges associated with this diversity measure, making it easier to leverage its properties.
@@ 1.1 Our Contributions
  • In Section 2, we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a warm-up, demonstrating how the competing terms in the objective function interact.
  • In Section 3, we present the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm and returning the best solution. In the special case of linear utility functions $g(S)=\sum_{v\in S}w(v)$, we prove that GIST offers a stronger $\nicefrac{{2}}{{3}}$-approximation guarantee.
[…]
  • In Section 5, our experiments show that GIST outperforms baselines for MDMS on synthetic data (in particular the classic greedy algorithm). Then we show that GIST can be used to build better single-shot subsets of training data for an image classification benchmark on ImageNet compared to margin sampling and $k$-center algorithms, demonstrating the benefit of optimizing for a blend of utility and diversity.
1.2 Related Work
@@
A closely related problem is $k$-center, which aims to find a subset of $k$ centers that minimize the largest distance of any point to its nearest center in $S$. Concretely, the objective is $\max_{v\in V}\min_{c\in S}\textnormal{dist}(v,c)$, and a greedy $2$-approximation algorithm is widely used in active learning applications (Sener and Savarese, 2018). The main difference between MDMS and $k$-center is the ability to benefit from the utility term, e.g., uncertainty values in an active learning setting. The second key difference is that we relax the requirement to cover every point in the dataset, which can be helpful for large datasets where we want to avoid instability due to certain corner or outlier points.
Bhaskara et al. (2016) study various distance-based diversity metrics subject to a cardinality constraint. In particular, they give a $\nicefrac{{1}}{{8}}$-approximation algorithm for the sum-min diversity objective defined as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ by rounding a linear programming relaxation. They complement this result by proving a conditional $\nicefrac{{1}}{{2}}$-hardness of approximation under the planted clique assumption. They also study the min-min objective $\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$, but they enforce a strict cardinality constraint $|S|=k$; otherwise, it is optimal to only output one or two points. Meinl et al. (2011) study the same objective subject to $|S|=k$ under the name $p$-dispersion (also called max-min or remote-edge) diversity maximization. This exact cardinality issue does not arise in MDMS because we need to balance the trade-off between competing monotone submodular utility and diversity terms.
Moumoulidou et al. (2020), Addanki et al. (2022), and Mahabadi and Trajanovski (2023) study diversity maximization under fairness constraints. Given a partitioning of the dataset, they ensure that a specified number of points $k_{i}$ are selected from each group $i\in[m]$. This fairness constraint considerably differentiates their problem setting from ours (i.e., $|S|=k=\sum_{i=1}^{m}k_{i}$). Moumoulidou et al. (2020); Addanki et al. (2022) maximize $\texttt{div}(S)$ in this setting. Mahabadi and Trajanovski (2023) consider sum-of-all pairwise distances $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ and sum-min objectives. However, none of these works add a utility term to the objective.
Borodin et al. (2017) optimize the combination of a submodular utility and sum-of-all pairwise distances in the selected set, making it the most similar work to MDMS. Unlike our $\texttt{div}(S)$ term, max-sum diversity is not very sensitive to duplicate points. In fact, adding $m$ copies of the same point to a set raises the increment of the diversity term by a factor of $m$ in their model. By using $\texttt{div}(S)$ instead, we can strictly avoid duplication across all selected points.
@@
In the realm of dataset curation and active learning, there have been several lines of work that study the combination of utility and diversity terms. Ash et al. (2020) invoke $k$-means++ seeding over the gradient space of a model in order to balance uncertainty and diversity. Wei et al. (2015) introduce several submodular objectives, e.g., facility location, and use them to diversify a set of uncertain examples in each active learning iteration. Citovsky et al. (2021) cluster examples that are represented by embeddings extracted from the penultimate layer of a partially trained DNN, and use these clusters to diversify uncertain examples in each iteration. Our work differs from these and several others (e.g., Kirsch et al. (2019); Zhdanov (2019)) in that we directly incorporate the utility and diversity terms in our objective function.
@@ 2 Preliminaries
Here we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a starting point.
Submodular set function.
For a set function $g:2^{V}\rightarrow\mathbb{R}$ and any $S,T\subseteq V$, let $g(T\mid S)=g(S\cup T)-g(T)$ denote the marginal gain of $g$ at $T$ with respect to $S$. A function $g$ is submodular if for every $S\subseteq T\subseteq V$ and $v\in V\setminus T$, we have $g(v\mid S)\geq g(v\mid T)$, where we overload the marginal gain notation for singleton sets. A submodular function $g$ is monotone if for every $S\subseteq T\subseteq V$, $g(S)\leq g(T)$.
[…]
Min-distance diversity.
For a set of $n$ points $V$ in a metric space, define the min-distance diversity function
@@
For $\texttt{div}(S)$ to be well-defined, we set the diversity to be the diameter of $V$ if $|S|\leq 1$, i.e., $d_{\max}=\max_{u,v\in V}\textnormal{dist}(u,v)$. This implies that it is always beneficial to select at least two points to maximize the overall objective.
We extend the distance function to take subsets $S\subseteq V$ as input, i.e., $\textnormal{dist}(u,S)=\min_{v\in S}\textnormal{dist}(u,v)$. For the sake of completeness, we define $\textnormal{dist}(u,\varnothing)=\infty$.
Optimization problem.
Define the objective function
$f(S)=g(S)+\texttt{div}(S).$ (2)
The goal of the min-distance diversification with monotone submodular utility (MDMS) problem is to maximize $f(S)$ subject to a cardinality constraint $k$:
@@
Remark 2.1.
The objective function $f(S)$ is not submodular. We present a counterexample in Appendix A.
Intersection graph.
Let $G_{d}(V)$ be the intersection graph of $V$ for distance threshold $d$, i.e., with nodes $V$ and edges
$E=\{(u,v)\in V^{2}:\textnormal{dist}(u,v)<d\}.$
Note that for any independent set $S$ of $G_{d}(V)$ and pair of distinct nodes $u,v\in S$, we have $\textnormal{dist}(u,v)\geq d$.
2.1 Warm-up: Simple $0.387$-Approximation Algorithm
The objective function $f(S)$ has two terms, each of which is easy to (approximately) optimize in isolation. For the $g(S)$ term, run the greedy algorithm to get $S_{1}$ satisfying
$g(S_{1})$ $\geq(1-\nicefrac{{1}}{{e}})\cdot\max_{|S|\leq k}g(S)$
$\geq(1-\nicefrac{{1}}{{e}})\cdot g(S^{*}).$
For the $\texttt{div}(S)$ term, let $S_{2}$ be a pair of diametrical points in $V$. The algorithm then flips a biased coin and outputs $S_{1}$ with probability $p\approx 0.613$, and $S_{2}$ otherwise. This achieves a $0.387$-approximation guarantee:
$\mathbb{E}[\textnormal{ALG}_{\textnormal{simple}}]$ $\geq p\cdot g(S_{1})+(1-p)\cdot\texttt{div}(S_{2})$
$\geq p\cdot(1-\nicefrac{{1}}{{e}})\cdot g(S^{*})+(1-p)\cdot\texttt{div}(S^{*})$
$=\frac{e-1}{2e-1}\cdot\textnormal{OPT},$
where we take $p=e/(2e-1)$ by solving $p\cdot(1-\nicefrac{{1}}{{e}})=1-p$. One of the main goals of this paper is to improve over this baseline and prove complementary hardness results.
@@ 3 Algorithm
In this section, we present the GIST algorithm and prove that it achieves an approximation ratio of $(\nicefrac{{1}}{{2}}-\varepsilon)$ for MDMS. GIST works by sweeping over multiple distance thresholds $d$ and calling a GreedyIndependentSet subroutine to find a high-value independent set of the intersection graph $G_{d}(V)$. This is in contrast to the simple $0.387$-approximation algorithm that just tries two extreme thresholds $D=\{0,d_{\max}\}$.
The GreedyIndependentSet heuristic builds a set $S$ (starting from the empty set) by repeatedly adding $v\in V\setminus S$ with the highest marginal gain with respect to the submodular function $g$ while satisfying $\textnormal{dist}(v,S)\geq d$. This subroutine runs until either $|S|=k$ or $S$ is a maximal independent set of $G_{d}(V)$.
@@
Algorithm 1 Diverse data summarization for submodular utilities using greedy-weighted maximal independent sets.
1:function GIST(points $V$, monotone submodular function $g:V\rightarrow\mathbb{R}_{\geq 0}$, budget $k$, error $\varepsilon$)
2:     Initialize $S\leftarrow\textnormal{{GreedyIndependentSet}}(V,g,0,k)$triangleright runs the classic greedy algorithm
@@
9:         Set $T\leftarrow\textnormal{{GreedyIndependentSet}}(V,g,d,k)$
@@
1:function GreedyIndependentSet(points $V$, monotone submodular function $g:V\rightarrow\mathbb{R}_{\geq 0}$, distance $d$, budget $k$)
@@
3:     for $i=1$ to $k$ do
4:         Let $C\leftarrow\{v\in V\setminus S:\textnormal{dist}(v,S)\geq d\}$
5:         if $C=\varnothing$ then
6:              return $S$triangleright $S$ is a maximal independent set of $G_{d}(V)$
7:         Find $t\leftarrow\operatorname*{arg\,max}_{v\in C}g(v\mid S)$
8:         Update $S\leftarrow S\cup\{t\}$      return $S$
@@
For any $\varepsilon>0$, the GIST algorithm outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{1}}{{2}}-\varepsilon)\cdot\textnormal{OPT}$.
The main building block in our design and analysis of GIST is the following lemma, which is inspired by the greedy $2$-approximation algorithm for metric $k$-centers and adapted for monotone submodular rewards.